# Harnessing natural and mechanical airflows for surface-based atmospheric pollutant removal Centre for Climate Repair, Cambridge, UK

### v4

@@ Results
Having established the atmospheric pollutant transport rates to the surfaces of various natural and mechanical environments, we assume ideal removal efficiencies and compare the potential atmospheric pollutant removal rates with the 1 GtCO2e/y target (20–100-year GWP) [20]. This target appears feasible for cities, solar farms, HVAC systems, filters, aeroplanes, trains and automobiles, based on the potential atmospheric pollutant flow rates to existing surfaces (Fig. 3). When focusing only on potential atmospheric pollutant flow rates to annually produced surfaces (Supplementary Figs. 2), feasible applications narrow to cities, solar farms, HVAC systems and filters.
@@
However, the removal efficiency is often suboptimal, falling short of the transport flux to surfaces. For example, DAC systems exhibit removal efficiencies around [46], such that the average pollutant removal potential is 1 GtCO2/y for HVAC systems. With 10 DAC systems removing, on average, around tCO2/y each [28], atmospheric removal rates are currently limited to approximately GtCO2/y (Fig. 3a). CH4 and N2O present additional challenges due to their low atmospheric concentrations but can also achieve removal efficiencies around [50]. However, most efficiency values are reported at elevated temperatures, with room temperature conversion rates still being an area of active research. When multiplied by the removal efficiency, cities, solar farms and HVAC systems have maximum pollutant removal potentials (20, 0.5, 2 GtCH4/y respectively), which are greater than atmospheric oxidation (0.3 GtCH4/y), providing a pathway that could surpass current emissions [9, 49]. PM2.5 is captured by HVAC filters which achieve removal efficiencies of around [42], such that the average pollutant removal potential is 0.01 GtPM2.5/y for HVAC filters. Atmospheric removal rates approach GtPM2.5/y (Fig. 3c), with a mean estimate of 100 million filters processing 100 m3/h at 99% efficiency [24].
Cost analyses
Finally, we estimate the potential cost of scaling CO2-sorption and CH4-catalyst technologies, highlighting their capacity for atmospheric pollutant removal across natural and mechanical airflows. These surface-based technologies are selected as promising approaches for GHG removal, with estimated costs ranging from $17 to $73 per m2 for CO2 sorbents and $14 to $60 per m2 for catalysts targeting CH4 based on [39, 29, 36, 18]. By contrast, atmospheric PM2.5, as well as the oxidised forms of NO, are captured when they interact with surfaces through deposition. As shown in Fig. 3c–d, dry deposition rates of GtPM2.5/y and GtNO/y align with atmospheric pollutant flow rate estimates to urban surfaces [53, 10], contributing around 1–10% of global deposition.
To evaluate the potential cost-efficiency of each application, we compare relative to atmospheric pollutant removal costs of $100 per tCO2e removed (20-year GWP), consistent with recommendations for GHG removal technologies [20]. We assume a removal efficiency of [46, 50, 32, 24], reflecting a realistic and optimistic average across surface-based removal technologies and pollutants. By dividing the estimated annual surface-based technology cost by the potential atmospheric pollutant removal rate, we derive the cost per tonne of CO2e removed for each application in Tab. 1. Estimates across natural and internal environments exceed the $100 per tCO2e removed target. However, HVAC filters achieve potential costs as low as $300 per tCO2 removed for sorption and $600 per tCO2e removed for catalyst technologies. This lower cost for HVAC filters is attributed to the high atmospheric pollutant flux to their surface (e.g., an average of mol/m2s for CO2, Fig. 2c), which reduces the influence of surface area on the potential removal cost.
We explore the potential of different catalysts for atmospheric methane removal by varying material properties and surface coverage. Palladium (Pd) and magnesium (Mg) are considered, spanning material costs from $10 to $104 per kg and densities from 1738 to 12023 kg/m3. We assume a catalyst particle size of 10-6m with 10% coverage. The cost per catalyst surface area is determined by multiplying the cost per kg by the catalyst mass and dividing by the catalyst surface area, yielding $0.02 to $120.2 per m2 of catalyst. Fluxes are computed for the total surface area using Tab. 1 and adjusted for catalyst coverage. Therefore, we estimate the cost per ton of CO2e removed to range from $ to $ for cities and $3.2 to $ for HVAC filters, the latter meeting the $100 per tCO2e target. Further reductions are possible by increasing catalyst coverage or reducing particle size, potentially bringing city-scale applications within target costs.
@@ Discussion
The potential for removing low concentrations of atmospheric pollutants as a mitigation strategy for climate change and public health crises has yet to be fully explored. Surface-based removal technologies have been proposed as a scalable approach to achieving these reductions. For example, it has been shown that DAC may remove CO2 from ambient air at scale [28], catalytic coatings in urban areas can reduce the concentration of volatile organic compounds (VOCs) and NO [43], and HEPA filters can lower PM levels [23]. To evaluate whether wider deployment is warranted, we have developed a framework to assess the potential scale of surface-based atmospheric pollutant removal, leveraging pre-existing airflows in cities, HVAC systems and transportation.
@@
Surface-based pollutant removal technologies in cities, solar farms, HVAC systems and filters could achieve potential atmospheric pollutant removal rates exceeding 1 GtCO2e per year (20–100 year GWP) [20], with optimisation for higher surface fluxes or lower materials costs potentially reducing the minimum cost estimate to under $100 per tCO2e. Building upon the existing capabilities of HVAC systems to remove PM and VOCs [23], by integrating other surface-based pollutant removal technologies, could enhance their cost-effectiveness and scalability. A similar integration of technologies was suggested during the COVID-19 pandemic, when HVAC systems were adapted to mitigate airborne viruses, using technologies such as UV-C light and antimicrobial coatings [48, 51]. This demonstrated the adaptability of HVAC systems and their potential for multifunctional air quality management to address broader environmental challenges. Furthermore, HVAC filters are regularly replaced, potentially making the maintenance of any surface-based removal technologies straightforward and allowing their performance to be sustained over time.
Refining pollutant removal models can not only enhance the accuracy of predictions but also provide insights into optimisation. Adsorption/desorption models, such as Langmuir–Hinshelwood or Eley–Rideal, and multicomponent diffusion models, such as Maxwell–Stefan, can help identify conditions that maximise reaction rates [37, 4]. Similarly, dispersion models, including Gaussian plume, Eulerian and Lagrangian approaches, can enable placement of removal surfaces to enhance the removal efficiency [6]. Lastly, our analyses excluded energy consumption to focus on existing natural and mechanical airflows, assuming thin and smooth coatings with consistent flow resistance. However, rough coatings with increased surface area could enhance turbulence, flow rates to the surface and therefore removal efficiency.
Scaling surface-based atmospheric pollutant removal technologies presents challenges that must be addressed, including material degradation, surface contamination and infrastructure integration. While sorption systems generally maintain consistent pollutant removal efficiencies [46, 42], the catalytic removal of CH4 is variable [50, 32] and may decrease over time or under varying conditions (e.g., relative humidity). Advances in materials and reaction mechanisms can improve these efficiencies and expand the applicability of catalysis. However, large-scale deployment of catalysts may introduce risks, such as the conversion of other species (e.g., VOCs) to more toxic forms.
@@
Methods
Scaling theory
To characterise atmospheric pollutant transport dynamics across various systems, we define velocity (, m/s), length (, m) and surface area (, m2) scales specific to each flow application (e.g., cities, HVAC systems, transport). We also estimate the number of occurrences () for each flow type to ensure realistic scaling for global estimations. The limitations of this scaling approach and additional context are provided in Supplementary Sec. 3.1.
The thickness of the turbulent () boundary layer (, m) is evaluated using , where the Reynolds number [41] and is the kinematic viscosity (m2/s). The normal pollutant flux (mol/m2/s) at the surface is determined by the diffusive flux, , as the normal velocity at the surface is zero. The normal pollutant flux is simplified here using Fick’s law, which assumes diffusion driven solely by concentration gradients without accounting for molecular interactions or multicomponent effects. This normal pollutant flux depends on the diffusivity (m2/s), concentration (mol/m3) and the boundary layer thickness (m) [13], where the Schmidt number . An effective diffusivity (m2/s) is included to account for enhanced mixing, which alters pollutant transport rates [34].
The formulas in (1)–(2) provide a means to estimate atmospheric pollutant transport through, and to the surface of, global environments. As shown in Supplementary Fig. 1, we calculate the global streamwise flow rate through the boundary layers of the world’s cities (solar farms) using [17, 44, 11].
Empirical relationships
Empirical models can be used to link flow characteristics with the potential for pollutant removal [13]. These relationships have been applied in CH4 and N2O removal studies via photocatalysis [50, 36]. These relationships are evaluated using , and , supplemented with measured values such as wall shear stresses ( in kg/ms2), mass transfer coefficients ( in m/s) and heat transfer coefficients ( in W/m2K). A detailed discussion of these empirical relationships and their applications can be found in Supplementary Sec. 3.2.
Analogies exist between key dimensionless numbers: the Nusselt number (), the Sherwood number () and the drag coefficient (), where is the thermal conductivity (W/mK) and is the fluid density (kg/m3) [13]. These analogies allow for the formulation of the normal pollutant flow rate to the surface of each application as follows
Following Tsopelakou et al. [50], we can use (3) to estimate the streamwise flow rate through each application as
where is the wetted perimeter and is the cross-sectional area of the channel or boundary layer.
Fully-developed velocity and pollutant profiles
We employ established formulas to determine fully-developed velocity and pollutant profiles, including the Monin–Obukhov (M-O) similarity theory for natural airflows and approximate solutions of the Navier-Stokes equations for internal airflows, with explanations provided in Supplementary Sec. 3.3.
For external airflows over cities or solar farms, we utilise M-O similarity theory [41], which incorporates the effects of buoyancy and surface roughness under varying thermal conditions. The normal flow rate of pollutants to the surface of the environment is
where represents the average disturbance height and is the roughness length. For , where is the viscous sublayer thickness, we recover the log-law profile. This log law applies to external airflows over smooth or mildly rough surfaces and can be used to calculate pollutant fluxes near vehicles [34].
For internal airflows within HVAC systems, we approximate solutions of the Navier-Stokes equations and boundary conditions [52]. The normal flow rate of pollutants to the channel walls is given by
where is the channel diameter. For internal airflows through HVAC filters, we adjust the velocity from the duct flow (with velocity , diameter and cross-channel area ) to the pleat flow (with velocity , diameter , filter sheet area , number ). We apply the principle of mass conservation to match average velocities, leading to . The flow rate of pollutants to the pleat channel walls is
The flow rate of pollutants to the fibre sheet is
where is the collection efficiency, related to the permeability and porosity of the filter medium [19].
Energy, power and drag measurements
We now outline formulas to determine the flow rate based on energy, power and drag measurements. More details are given in Supplementary Sec. 3.4.
For external flows over urban areas or solar farms, we establish a control volume with a length and height . In the case of uni-directional, steady-state flow over a surface with a heat flux (W/m2), energy conservation dictates that the energy entering the control volume must balance with the energy exiting it [13]. This relationship can be expressed as
where (m2/s2K) is the specific heat capacity of air at constant pressure and (K) is the temperature difference between the surface and the surrounding atmosphere. The flow rates are related using the scaling theory discussed in Supplementary Sec. 3.1, such that .
where is the leakage factor [40].
For external airflows acting on transportation systems, the streamwise flow rate through the boundary layer can be estimated using the drag force acting on the aeroplane, train or automobile [52]. This is expressed by the equation
where (N) is the drag force.
Industry standards
For internal airflows through HVAC ducting, the streamwise flow rate can be evaluated using three estimates [1]. First, can be expressed in terms of the cubic feet per minute per person (, ft3/min), given by
where (m3) is the volume of air within the building.

### v5

@@ Results
CO2 using sorption
Having established the flow rate of CO2 to the surfaces of various natural and mechanical environments, we first assess the removal potential by multiplying that flow rate by an optimal removal efficiency for sorbent technologies, , comparing the potential removal rate of CO2 with the 1 GtCO2/y target 21. Our analysis suggests that potential removal rates exceeding 1 GtCO2/y using sorbents appear feasible for cities (based on the IQR: to GtCO2/y), solar farms ( to GtCO2/y), HVAC systems ( to GtCO2/y), filters ( to GtCO2/y), aeroplanes ( to GtCO2/y), trains ( to GtCO2/y) and automobiles ( to GtCO2/y), based on flow rates of CO2 to all existing surfaces (Fig. 3A). As all existing surfaces are assumed to be coated with sorbent technologies, the potential removal rates represent a theoretical upper bound. If one only considers flow rates to annually produced surfaces (Fig. S2A), which can provide an estimate for potential deployment rate or market penetration, feasible applications narrow to cities, solar farms, HVAC systems and filters. For example, if sorbents were applied only to new building materials, the transition from annually produced surfaces to total existing surfaces would occur gradually over decades (Tab. S2).
However, the feasibility of the above results may be uncertain, because the removal efficiency is often suboptimal, such that removal falls short of the flow rate to surfaces. For example, DAC systems using sorbents exhibit removal efficiencies of around (based on Soeherman et al. 52, which reviews various CO2-removal technologies), giving a median removal potential of GtCO2/y for cities, GtCO2/y for HVAC systems and GtCO2/y for automobiles with sorbent surfaces. We can compare these potential removal rates with current DAC systems which are removing around tCO2/y (based on McQueen et al. 32, which reviews DAC approaches in the US) or GtCO2/y (Fig. 3A).
CH4 using catalysis
Methane removal is more challenging due to its lower atmospheric concentration, but catalytic approaches can reach removal efficiencies of around (based on Tsopelakou et al. 56, which reviews CH4 photocatalyst data). However, most removal efficiencies are measured at elevated temperatures, and achieving comparable removal rates at room temperature remains an active area of research. When the removal efficiency is multiplied by the flow rate of CH4 to their surfaces, cities and HVAC systems have maximum removal potentials (upper 5% of of estimated values) of 10 and 1 GtCH4/y respectively (Fig. 3B), exceeding estimates of atmospheric oxidation (0.3 GtCH4/y) and providing a pathway that could surpass current CH4 emissions (based on Dlugokencky et al. 9, which constrains global CH4 emissions and sinks). These estimates assume complete surface coverage of each application with catalyst and sustained catalytic activity, and should be interpreted as theoretical upper bounds. Real-world constraints have varied pollutant reductions in field demonstrations of photocatalytic surfaces 44, 30, which operate at much smaller scales than the global surfaces considered here.
PM2.5 using filtration and oxidised forms of NO via deposition
PM2.5 is captured by HVAC filters which achieve removal efficiencies of around (based on Schroth 48, which compares various filter classes), such that the average removal potential is GtPM2.5/y (Fig. 3C). Atmospheric removal rates approach GtPM2.5/y (Fig. 3C), with estimates of 100 million filters processing 100 m3/h at 99% removal efficiency (based on Lowther et al. 27, which compares PM removal across rooms in China). These removal efficiencies represent realistic operational averages that depend on filter replacement cycles and are higher than those of sorption or catalyst technologies. By contrast, PM2.5, as well as the oxidised forms of NO, are captured when they interact with surfaces through deposition. As shown in Fig. 3C–D, dry deposition rates of GtPM2.5/y and GtNO/y align with atmospheric pollutant flow rate estimates to urban surfaces (based on Wu et al.59 which measured dry and wet deposition of particulate matter during summer in Beijing and Dore et al.10 which modelled the atmospheric transport and deposition of sulphur and nitrogen across the UK), contributing around 1–10% of global deposition.
Materials costs
CO2 using sorption and CH4 using catalysis via literature estimates
Finally, we estimate the potential materials cost of scaling CO2-sorption and CH4-catalyst technologies, highlighting their capacity for atmospheric removal if deployed in different environments. These technologies are selected as promising approaches for GHG removal, with reported material-only costs ranging from $16 to $50 per m2 for CO2 sorbents (based on values given in Fig. 28 of Sanz et al. 45 and Fig. 7 of Kulkarni et al. 24) and $15 to $43 per m2 for CH4 catalysts (based on values given in the SI of Randall et al. 42 and Hickey et al. 18). These values are estimates based on material prices and are combined with our calculated surface areas for different environments. They mostly exclude fabrication, installation, maintenance and operational energy inputs (e.g., UV illumination for photocatalysts), all of which could increase costs. Installed costs will vary with the application, and a full cost analysis would be required for each application.
To evaluate cost-efficiency, we compare these materials-only cost estimates with a benchmark of $100 per tCO2e removed (20-year GWP), in line with recommended targets for GHG-removal technologies 21. The rationale for this is that if the materials cost alone approaches or exceeds the targeted installed cost, this should be flagged, as it may indicate that further investigation of this material has limited value. We assume a removal efficiency of for CO2 sorbents 52 and for CH4 catalysts 56. For each application in Tab. 1, we calculate the materials cost per tonne of CO2e removed by dividing the estimated material cost of sorbent or catalyst technologies by the corresponding CO2 or CH4 removal rate. As expected, most estimates across natural and internal environments exceed the $100 per tCO2e removed target. However, HVAC filters achieve potential materials costs as low as $600 per tCO2 removed (range: $ to per tCO2 removed) for sorption, and as low as $2000 per tCO2e removed ($ to per tCO2e removed) for CH4-catalyst technologies (Tab. 1). This relatively lower materials cost arises from the high pollutant flux to HVAC filter surfaces (e.g., an average of mol/m2s for CO2, Fig. 2C), which reduces the surface area requirement per unit of pollutant removal.
@@
CH4 using catalysis via modelling
We explore the potential of different catalysts for atmospheric methane removal by varying material properties and surface coverage. Palladium (Pd) and magnesium (Mg) are considered, spanning material costs from $10 to $104 per kg and densities from 1738 to 12023 kg/m3. We assume a catalyst particle size of 10-6 m with 10% coverage. The materials cost per catalyst surface area is determined by multiplying the materials cost per kg by the catalyst mass and dividing by the catalyst surface area, yielding $0.02 to $120.2 per m2 of catalyst. Fluxes are computed for the total surface area using Table 1 and adjusted for catalyst coverage. Therefore, we estimate the materials cost per ton of CO2e removed to range from $ to $ for cities and $ to $ for HVAC filters, the latter below the $100 per tCO2e target. This modelling framework suggests that further cost reductions are possible by increasing catalyst coverage or reducing particle size, potentially bringing city-scale applications within target costs. While consistent with lab-scale observations 56, this has not been experimentally validated in city-scale applications. Furthermore, reducing particle size may increase processing and handling costs (not included here), so optimisation must balance activity gains with fabrication and deployment expenses. Realistic deployment in specific applications will therefore require a complete cost evaluation. The values presented here are best interpreted as material-cost estimates based on potential surface area, allowing us to compare across applications, with full system costs expected to be higher once fabrication, integration and maintenance are included.
@@ Discussion
The potential for removing low concentrations of atmospheric pollutants, as a mitigation strategy for climate change and public-health crises, has yet to be fully explored. Surface-based removal technologies have been proposed as a potentially scalable approach, with examples including sorbents for CO2 capture 32, catalysts to reduce VOCs and NO 49, and HEPA filters for PM removal 26. Recent field deployments show photocatalytic systems achieving NO removal of up to 66% at street level and 82% on coated wall surfaces, while in enclosed spaces, O3 reductions of up to 23% have been observed 44. Further, Ma et al. 30 discusses the potential of photocatalytic systems to deliver atmospheric pollutant reductions outside the laboratory in cities, for species including NO, VOCs and O3. Building on this evidence, we quantify global removal potential from pollutant fluxes, including high-GWP gases such as CH4, thereby linking air-quality benefits with climate-relevant mitigation. To achieve this, we developed a framework that leverages pre-existing airflows and surfaces across cities, HVAC systems and transportation, while recognising that other environments, such as cooling towers or clothing, could also potentially contribute to pollutant removal.
@@
Surface-based pollutant removal technologies in cities, solar farms, HVAC systems and filters could achieve potential atmospheric pollutant removal rates exceeding 1 GtCO2e per year (20 year GWP) 21, with optimisation for higher surface fluxes or cheaper materials potentially reducing the minimum materials cost estimate to below $100 per tCO2e. These potential removal rates represent theoretical upper bounds, assuming all existing surfaces are coated with removal technologies at literature-based lab efficiencies, and are intended to enable comparison across applications rather than specific deployment. To approximate potential deployment rates, we also evaluate potential pollutant removal rates for annually produced surfaces coated with removal technologies (Fig. S2). Building upon the existing capabilities of HVAC systems to remove PM and VOCs 26, by integrating other surface-based pollutant removal technologies, could enhance their cost-effectiveness and scalability. A similar integration of technologies was suggested during the COVID-19 pandemic, when HVAC systems were adapted to mitigate airborne viruses, using technologies such as UV-C light and antimicrobial coatings 54, 57. This demonstrated the adaptability of HVAC systems and their potential for multifunctional air quality management to address broader environmental challenges. Furthermore, HVAC filters are regularly replaced, potentially making the maintenance of any surface-based removal technologies straightforward and allowing their performance to be sustained over time.
For context, the median potential removal rate of CO2 with sorbents on city surfaces (30 GtCO2/y; Fig. 3A) corresponds to nearly three-quarters of global anthropogenic CO2 emissions rates (40 Gt/y) in Friedlingstein et al. 14. For comparison with mitigation needs, the IPCC Sixth Assessment Report projects that achieving net-zero CO2 by 2050 requires removal rates of up to 18 GtCO2/year, depending on the pathway 20. Our median potential removal rate of CO2 with sorbents on city surfaces therefore indicates an application which could be targeted to achieve net-zero. For CH4, the maximum modelled potential removal rate with catalysis on city surfaces (10 GtCH4/y) exceeds current atmospheric oxidation rates (0.3 GtCH4/y) 9, the latter being approximately comparable to global CH4 emissions rates. Because the atmospheric lifetime of CH4 is around 10 years 4, even partial realisation of removal at the scales indicated here could lead to measurable decreases in global mean temperature within decades, consistent with near-term climate goals 20. For NO, projected removal rates with catalysis on city surfaces are around 10% of global deposition rates, estimated using Zhao et al. 64 (Fig. 3D), while for PM2.5, the flow rate to city surfaces is approximately 1% global dry deposition rates, estimated using Wu et al. 59 (Fig. 3C). However, removal efficiencies in application may be lower due to suboptimal operating conditions, material degradation or limited surface coverage, and reaching significant atmospheric impact would require sustained deployment over multiple years.
Although our main analysis focused on CO2, CH4, PM2.5 and NO, the methodology is general and can be applied to many atmospheric pollutants and removal technologies. One must specify the pollutant’s atmospheric concentration, diffusivity, surface area and airflow of the environment, calculating the atmospheric flow rate using Methods and applying the removal efficiency of the chosen technology. For example, O3 is a short-lived climate pollutant linked to respiratory illness and crop yield losses 63, with typical atmospheric concentrations of 60 ppb ( mol/m3) and diffusivity of m2/s. Using these values in (3) for city environments (Tab. S2), gives an atmospheric flow rate to all city surfaces of around mol/s. With a photocatalytic removal efficiency of 0.5 given by Ohtani et al. 34 for TiO2, the above flow rate corresponds to a removal potential of GtO3/year, which represents several percent of the atmospheric O3 burden.
Scaling surface-based atmospheric pollutant removal technologies presents several challenges that must be addressed in future research, including material degradation, surface contamination and infrastructure integration. Sorption systems generally maintain predictable CO2 removal efficiencies that decline as sorbent materials become saturated 52, whereas the catalytic removal of CH4 is more variable 56, 37, and can fluctuate under varying conditions (e.g., relative humidity, light intensity). Advances in materials and reaction mechanisms can enhance these removal efficiencies and broaden the applicability of catalysis. However, large-scale deployment of catalysts may result in undesirable reactions, such as the partial oxidation of CH4 yielding formaldehyde 38, or the reduction of NO producing N2O 23. The distribution of such products can vary depending on catalyst composition, morphology, humidity and pollutant mixtures. Mitigation measures include tailoring catalysts to favour complete oxidation of CH4 to CO2 and H2O while suppressing partial oxidation pathways 38, immobilising catalysts on durable supports to minimise particle release 65 and operating under controlled temperature–humidity conditions to reduce N2O formation from NO reduction 23. In HVAC systems, catalytic removal could be combined with downstream filtration to capture any harmful intermediates before air is recirculated. Additional risks, such as catalyst poisoning by sulphur or siloxane species 66, can be mitigated via upstream pre-filtration and periodic regeneration of catalysts. Incorporating these safeguards into deployment strategies may enable the benefits of large-scale catalysis to be realised while protecting environmental and public health.
@@
We acknowledge Grantham Foundation for supporting this research.
Additional supplementary Figures S1–S2 and Tables S1–S5 are provided in the Supporting Information.